Edit View History Bookmarks Tools Window Help 72%..) Monu Do d 48421 884 2018 Onine Stat 2104 Select Toplos In StatL Bus Analytle Cristian Smith &16/11/18 10:12 PM Homework: HW#2: CHAPTERS 7 & 8 M) Score: 0 of 1 p 15 of 22 (14 completel MN Score: 5838%, 12.4 of 22 pts 7.6.75-T What are the hypotheses for this est? p-06 p#061 Caloulae the value of the z-atatiotic for this es Round so beo decimal places as neoded
Answer #1

Ans:

Sample proportion=(375+62)/756=0.578

Test statistic:

z=(0.578-0.61)/sqrt(0.61*(1-0.61)/756)

z=-1.80

p-value=P(z<-1.80)=0.0359

(without rounding off z value,p-value=0.0358)

As,p-value>0.01,we fail to reject the null hypothesis.